# Marharyta Domnich, Julius Välja, Rasmus Moorits Veski, Giacomo Magnifico, Kadi Tulver, Eduard Barbu, Raul Vicente

### v1

@@ Introduction
Different people’s preferences for explanations exhibit significant variability. To explore the effects of this, an experiment was carried out with a dataset based on specific participants’ answers, instead of the sample averages. To ensure that these participants represent different subgroups of participants, t-SNE was used to reduce the dimensionality of the data and DBSCAN clustering to cluster the results. The goal of clustering was to discern the largest clusters present in the data, after which a random participant was chosen from each of the four largest clusters. The results of clustering and participant selection can be viewed in Appendix D, Figure D.1.
@@
The results of this process were varied with accuracies ranging from 58% to 69% for zero-shot. Table 5 shows that the LLM’s predictions improved significantly after fine-tuning, reaching accuracies over all metrics of $\sim\!\!90\%$ for 3 participants. One participant appeared to be less consistent, as the model managed to simulate their answers with an accuracy of only 66%. This leads to two conclusions: while LLM’s biases and preferences can be tuned to match specific participants to a great extent, some participants’ preferences prove significantly more difficult to mimic. However, since this comparison only contained 4 participants and 30 explanations, these conclusions should be considered tentative.
@@ Discussion
The traditional assessment of counterfactual explanations often overlooks human aspects, relying either on inconsistent quantitative metrics (frequently used both within objective function optimization and for evaluation (Cheng, Ming, and Qu 2021)) or on user studies that focus on a specific subset of individuals, lacking comparability over time and methods. To address this, we developed a novel dataset of counterfactual explanations, evaluated by human participants, which demonstrated a diverse spread of evaluations across all metrics, highlighting its applicability in different contexts. Utilizing this dataset to fine-tune LLMs demonstrated promising results, achieving an 85% accuracy, suggesting they can be used to approximate human judgment across various metrics. Furthermore, the zero-shot LLM performance was already notable, achieving up to 63% accuracy. Our experiments also indicate the potential to fine-tune models to individual experts, targeting specific knowledge or preferences.
@@
Despite the potential, it is crucial to acknowledge that LLMs do not replace the nuanced insights provided by human evaluations. Instead, they can serve as a complementary tool, enhancing scalability and reducing the resources required for broad assessments across multiple frameworks. Moreover, we propose exploring the idea of integrating this model within a human-in-the-loop approach to produce a hybrid model that could refine the quality of counterfactual explanations during the generation process (i.e. creating an LLM-in-the-loop instead of a human) (Abrate et al. 2024), leveraging the strengths of both automated and human evaluations.
@@ Acknowledgments
This research was supported by the Estonian Research Council Grants PRG1604, the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 952060 (Trust AI), by the Estonian Centre of Excellence in Artificial Intelligence (EXAI), by the Estonian Ministry of Education and Research.
@@ Appendix A Appendix A
Imagine you are in this scenario: ”You are a 31-year-old divorced woman. You have a high-school education and you work 20 hours per week.” Current outcome: You are earning less than the average salary. Useful context: the standard full-time workload is 40 hours per week. ”To earn more than the average salary, you would need to make the following changes: Increase your education level from high-school to Bachelor’s degree.”

### v2

@@ Introduction
Different people’s preferences for explanations exhibit significant variability. To explore the effects of this, we conducted an experiment using a dataset based on specific participants’ answers rather than sample averages. To ensure that these participants represent different subgroups of participants, t-SNE was used to reduce the dimensionality of the data, and DBSCAN clustering identified the largest clusters. random participant was selected from each of the four largest clusters. The results of clustering and participant selection can be viewed in Appendix D, Figure D.1.
@@
The results of this process varied, with accuracies ranging from 58% to 69% for zero-shot evaluation. Table 5 shows that the LLM’s predictions improved significantly after fine-tuning, reaching accuracies over all metrics of $\sim\!\!90\%$ for 3 participants. One participant appeared to be less consistent, as the model managed to simulate their answers with an accuracy of only 66%. This leads to two conclusions: while LLM’s biases and preferences can be tuned to match specific participants to a great extent, some participants’ preferences prove significantly more difficult to mimic. However, since this comparison only contained 4 participants and 30 explanations, these conclusions should be considered tentative.
@@ Discussion
The traditional assessment of counterfactual explanations often overlooks human aspects, relying either on inconsistent quantitative metrics (frequently used both within objective function optimization and for evaluation (Cheng, Ming, and Qu 2021)) or on user studies that focus on a specific subset of individuals, lacking comparability over time and methods. To address this, we developed a novel dataset of counterfactual explanations, evaluated by human participants, which demonstrated a diverse spread of evaluations across all metrics, highlighting its applicability in different contexts. Utilizing this dataset to fine-tune LLMs demonstrated promising results, achieving an 85% accuracy, suggesting they can be used to approximate human judgment across various metrics. Furthermore, the zero-shot LLM performance was already notable, achieving up to 63% accuracy. Our experiments also indicate the potential to fine-tune models to individual experts, to target specific expertise or individual preferences.
@@
Despite the potential, it is crucial to acknowledge that LLMs do not replace the nuanced insights provided by human evaluations. Instead, they can serve as a complementary tool, enhancing scalability and reducing the resources required for broad assessments across multiple frameworks. Moreover, we propose exploring the idea of integrating this model within a human-in-the-loop approach to produce a hybrid model that could refine the quality of counterfactual explanations during the generation process, effectively creating an LLM-in-the-loop approach instead of a human (Abrate et al. 2024), combining the strengths of automated and human evaluations.
@@ Acknowledgments
This research was supported by Estonian Research Council Grants PRG1604, the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No. 952060 (Trust AI), the Estonian Centre of Excellence in Artificial Intelligence (EXAI), the Estonian Ministry of Education and Research.
@@ Appendix A Appendix A
Imagine you are in this scenario: ”You are a 31-year-old person, who has had one pregnancy, and who has the following medical readings: glucose level: 173 mg/dL, diastolic blood pressure: 82 mm Hg, skin thickness: 48 mm, Insulin: 160 $\mu$IU/ml, body mass index (BMI): 32.8.” Current outcome: You have an increased risk of diabetes. Useful context: Normal glucose levels are 70-130, 140-200 is prediabetes, 200+ is diabetes. Healthy BMI is 18.5 - 25, 25-30 is considered overweight, 30+ is considered obese. Normal diastolic blood pressure is roughly below 80 mm Hg. Normal insulin is between 16 and 166 $\mu$IU/ml. ”To not be at increased risk of diabetes you would need to make the following changes: Decrease your glucose level from 173 to 130. And, increase your insulin level from 160 to 181.”
@@
Demographic Description Number Percentage (%)
Age Group 18-24 years old 64 31.07
25-34 years old 95 46.12
35-44 years old 25 12.14
45-54 years old 15 7.28
55-64 years old 7 3.40
Nationality South African 32 15.53
Polish 30 14.56
British 27 13.11
Mexican 20 9.71
Portuguese 19 9.22
Chilean 12 5.83
Italian 8 3.88
Greek 7 3.40
Hungarian 7 3.40
USA 6 2.91
Other 38 18.45
Education Level Primary school or lower 1 0.49
High school 69 33.50
Bachelor’s degree or equivalent 93 45.15
Master’s degree or equivalent 39 18.93
Doctoral degree or equivalent 4 1.94
English Proficiency Native speaker / Fully proficient 166 80.58
Moderately proficient 40 19.42
Experience in Machine Learning No experience 100 48.54
Some experience 89 43.20
I am studying in a related field 12 5.83
I work in the field / Extensive experience 5 2.43
Experience with Causality Frameworks No 145 70.39
I am familiar with the general concept 57 27.67
I have previous experience with them 4 1.94
Medical Background No 181 87.86
I am currently studying medicine or a related field 10 4.85
I work with medical data or in a field related to medicine 8 3.88
I have a degree in medicine or a related field 4 1.94
I work in the field of medicine 3 1.46
Table A.3: Demographics of the participants.